Consider the reaction C(s)+CO2(g)⇌2CO(g).

When 1.66 mol of CO2 and an excess of solid carbon are heated in a 21.2 L container at 1100K, the equilibrium concentration of CO is 7.19×10−2 M .

Part A What is the equilibrium concentration of CO2?

Part B What is the value of the equilibrium constant Kc at 1100 K?
Answer #1

in the equilibrium expression we dont want to tbother about the solide

convert moles in to concentration

M = moles / volume in liters = 1.66 mol / 21.2 L = 0.0755 mol

C(s)+ CO2(g) ⇌ 2CO(g)

I 0.0755 0

C -x +2x

E 0.0755 -x +2x

but in problem he has already given equilibrium concentration of CO = 7.19×10−2 M = 2x

x = 7.19×10−2 M / 2 = 3.595 x 10-2 M

equilibrium concentration of CO2 = 0.0755 - x = 0.0755 -7.19×10−2 M

= 0.0755 - 0.0719

= 0.0036 M

now Kc = [CO]2 / [CO2]

Kc = [7.19×10−2 ]2 / [0.0036]

Kc = 0.005169 / 0.0036

Kc = 1.436